question 51 of chapter 9 centre of mass and collision

Grade:11

1 Answers

we will so
6 years ago
If both the blocks are pulled by some force, they suddenly move with some acceleration and
instantaneously stop at same position where the elongation of spring is maximum.
 Let x1, x2  extension by block m1 and m2
Total work done = Fx1 + Fx2
 Increase the potential energy of spring = (1/2) K (x1+ x2)2 …(2)
Equating (1) and (2)
F(x1 + x2) = (1/2) K (x1+ x2)2  (x1+ x2) =
K
2F
Since the net external force on the two blocks is zero thus same force act on opposite direction.
 m1x1 = m2x2 …(3)
And (x1+ x2) =
K
2F